solve by completing the square
4x^2+2x-3=0

Answer by stanbon(75887)

4x^2+2x-3=0
Rewrite as follows:
4(x^2+(1/2)x+?) = 3+4*?
Complete the square on the left and keep the equation balanced.
4(x^2+(1/2)x+(1/4)^2) = 3 + 4*(1/4)^2
Factor the left and simplify the right.
4(x+(1/2))^2 = 3+1/4
Divide both sides by 4
(x+(1/2)^2 = 13/16
Take the square root of both sides.
(x+(1/2)) = [sqrt13]/4 or x+(1/2) = -[sqrt13]/4
Rearrange and solve for x:
x = -(1/2)+[sqrt13]/4 or x = -(1/2)-[sqrt13]/4
===================
Cheers,
Stan H.